RhondaSommer (rhondasommer):

Hey, can someone help me with a few trig/pre-calc questions? I'll fan you and I'll medal you on each one (create different questions for each one) . I just need help to go through it. Thanks :)

jimthompson5910 (jim_thompson5910):

Simply post your full question to get started. Please only post one question per thread to avoid lag and clutter

RhondaSommer (rhondasommer):

I will post in just a second. Im having a hard time with screenshot sorry

TheSmartOne (thesmartone):

It hass been 20 minutes since you posted the question you're still having a problem posting a screenshot. If you want to know how to post a screenshot, check out this: http://goo.gl/Xr1itf

RhondaSommer (rhondasommer):

sorry, my computer got wet and now it acts up from time to time. :P

jimthompson5910 (jim_thompson5910):

That's ok. Take all the time you need.

jimthompson5910 (jim_thompson5910):

Do you agree that \[\Large \sin\left(\frac{1}{2}x\right)\] is the same as \[\Large 1*\sin\left(\frac{1}{2}x-0\right)+0\] ??

RhondaSommer (rhondasommer):

yes. right? im pretty sure anyway. ha, I used to love math and then they added words to it and i got confused XD

jimthompson5910 (jim_thompson5910):

Notice how the color-coded terms match up \[\Large {\color{red}{1}}*\sin\left({\color{blue}{\frac{1}{2}}}x-{\color{green}{0}}\right)+{\color{purple}{0}}\] \[\Large {\color{red}{A}}*\sin\left({\color{blue}{B}}x-{\color{green}{C}}\right)+{\color{purple}{D}}\] where... |A| = amplitude 2pi/B = period C/B = phase shift y = D is the equation of the midline

RhondaSommer (rhondasommer):

thats right. its alwayts (2pi)/b isnt it?

jimthompson5910 (jim_thompson5910):

yes the period is 2pi/B in this case, B = 1/2 or you can think of it as B = 0.5 if fractions are a pain for you

RhondaSommer (rhondasommer):

so 1/2 pi

jimthompson5910 (jim_thompson5910):

Period \[\Large T = \frac{2\pi}{B}\] \[\Large T = \frac{2\pi}{0.5}\] \[\Large T = ???\] leave the answer in terms of pi

RhondaSommer (rhondasommer):

2/0.5 is 4 soooo.... 4pi?

jimthompson5910 (jim_thompson5910):

yep, you got it

RhondaSommer (rhondasommer):

cool well what if....okay ill ask you in the next question.